OpenStudy (anonymous):

PLZ HELP ASAP @aaronq @aum @phi @texaschic101 @cwrw238 @johnweldon1993 @jhonyy9 @mathstudent55 @dan815

OpenStudy (anonymous):

The table below shows the ticket rates for whale watching trips offered by Dolphin Tricks & Tours: Age Ticket Price Under three years free Three to 12 years $35 Over 12 years $46 On a certain day, the company took 88 people on whale watching trips. There were 44 children aged 12 and under, of which some children were under three years. If x represents the number of children under three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day?

OpenStudy (anonymous):

A system of equations is shown below: 6x - 5y = 5 3x + 5y = 4 The x-coordinate of the solution to this system of equations is _____.

OpenStudy (anonymous):

The table below shows two equations: Equation 1 |3x – 1| + 7 = 2 Equation 2 |2x + 1| + 4 = 3 Which statement is true about the solution to the two equations? Equation 1 and equation 2 have no solutions. Equation 1 has no solution, and equation 2 has solutions x = 0, 1. The solutions to equation 1 are x = -1.3, 2, and equation 2 has no solution. The solutions to equation 1 are x = -1.3, 2, and equation 2 has solutions x = 0, 1.

OpenStudy (cwrw238):

the absolute values (in the vertical lines) are POSITIVE so isolate these in both equations by adding 7 and 4 in equations 1 and 2 respectively and see what you get

OpenStudy (cwrw238):

* oops - sorry - subtract 7 and 4 - not add

OpenStudy (anonymous):

3

OpenStudy (cwrw238):

no solution to either equation